An advertising media analyst wants to estimate the mean weekly amount of time consumers’ spend watching television daily. Based on previous studies, the standard deviation is assumed to be 20 minutes. The media analyst wasn’t to estimate, with 99% confidence, the mean weekly amount of time to within +/- 5 minutes.
a) What sample sizes is needed
b) If 95% confidence is desired, how many consumers need to be selected?
Answer #1

DATE:29-07-2020

ANSWER:--------

a). for 99.7. Confidence interval, 2= 2.576. Hence Sample size needed. 2 n = (20 - (2.526x24] [SD-02 20 mg E = +5. (10.304)6). for 95% CI, 2=1.96 Hence, sample size needed. n- 2013 (1-96x20] = (7.84² = 61.46 ~ 62. Hence decrealeng ci level decrease